1)Twelve years ago, Mussab was five times as old as KhaledIn year's time Khaled will be half Mussab's age. How old is each man at the moment?


1
Expert's answer
2021-06-01T12:51:42-0400

Let KhaledIn is "x" years old, and Mussab is "y" years old. Then we have the following equations: "y-20=5(x-20)" and "y+1=2(x+1)." It follows that "y=5x-80" and "y=2x+1". Therefore, "5x-80=2x+1", and thus "3x=81." We conclude that "x=27" and "y=2\cdot 27+1=55."


Answer: KhaledIn is "27" years old, and Mussab is "55" years old at the moment.